### v1

# A Deep Reinforcement Learning-Based TCP Congestion Control Algorithm: Design, Simulation, and Evaluation

Abstract
This paper presents a novel TCP congestion control algorithm based on Deep Reinforcement Learning. The proposed approach utilizes Deep Q-Networks to optimize the congestion window (cWnd) by observing key network parameters and taking real-time actions. The algorithm is trained and evaluated within the NS-3 network simulator using the OpenGym interface. The results demonstrate significant improvements over traditional TCP New Reno in terms of latency and throughput, with better adaptability to changing network conditions. This study emphasizes the potential of reinforcement learning techniques for solving complex congestion control problems in modern networks.
Index Terms:
TCP Congestion Control, Reinforcement Learning, Deep Q-Network, NS-3, Network Simulation, AI in Networking
I Introduction
Transmission Control Protocol (TCP) is a fundamental component of modern computer networks, enabling reliable data transfer across the Internet. However, congestion can occur when network links reach their capacity limits due to excessive traffic, leading to increased latency, packet loss, and reduced throughput. Traditional TCP congestion control algorithms often fail to adapt effectively to dynamic network conditions, resulting in suboptimal performance. This study introduces a Deep Q-Network (DQN)-based congestion control algorithm designed to optimize the congestion window (cWnd) in real time based on network state observations. The algorithm is modeled using Markov Decision Processes (MDPs) and implemented using Python, C++, and TensorFlow. The model is evaluated using the NS-3 network simulator running on Linux via the Windows Subsystem for Linux (WSL). Simulation results indicate that the RL-based algorithm achieves 12.51% lower latency and a 68.31% higher data transmission rate compared to the conventional TCP New Reno algorithm. Furthermore, it demonstrates fast adaptation to variable network conditions. These results suggest that the proposed AI-based algorithm holds promise, particularly in mobile networks and Internet of Things (IoT) environments where network dynamics vary rapidly.
II Background and Related Work
Definitions:
Network Topology: Network topology refers to the physical or logical arrangements that define the interconnection of devices in a computer network and the logic of data exchange between them. Gains in reliability, performance, energy consumption, and complexity are achieved through network topology [1].
TCP: The purpose of TCP is to ensure secure data transfers in computer networks. Transmitted data reaches the destination completely, in order, and accurately. Data transfer first establishes a connection between the sender and receiver and manages it through a ”three-way handshake” process. If errors occur in the data during transmission, TCP detects and corrects the errors, and retransmits the data. Large data groups are divided into smaller groups and transmitted to the receiver as combined groups [2].
Congestion Window: Network congestion control has evolved significantly since the late 1980s. Jacobson [3] introduced the concept of ”packet conservation” and proposed solutions like slow start and congestion windows to prevent congestion-induced collapses. These innovations were implemented to improve TCP’s performance under congested conditions. Subsequent research has focused on specific performance issues. Cardwell et al. [4] presented BBR as a solution to buffer bloat and low throughput problems associated with loss-based congestion control in networks with large or small bottleneck buffers. The field has expanded to include wireless networks, with various TCP versions and protocols specifically developed for congestion avoidance and control in wireless environments [5]. This evolution reflects the increasing complexity and diversity of modern network infrastructures, indicating a shift in congestion control research from addressing network stability to overcoming specific performance challenges in different network types.
Network Simulator 3: With the growth of computer networks and the development of complex scenarios, network simulators have begun to play a significant role in the research area. Simulators are useful tools for implementing new security solutions, managing time and resources, and evaluating performance. By enabling the virtualization of the environment, they allow the analysis of network performance under different scenarios and their modeling according to specific criteria. NS3 offers features such as working compatibility, robust memory management, debugging, coding in C++, object-oriented concepts, and models supported by NS-2. It is highly suitable for wireless networks [6].
Deep Q-Learning: Deep Q-Learning (DQL) is a significant algorithm in reinforcement learning that uses Deep Q-Networks (DQN) to approximate the Q-function in sequential decision-making problems [7]. Although successful in laboratory conditions, its real-world application is limited due to gaps between theory and practice. DQL has been applied to various fields such as natural language processing [8], image processing, and financial forecasting [9]. The algorithm’s success has led to increased interest in understanding human learning processes and has advantages over other learning algorithms as it does not require prior data [7]. Recent advancements in GPU technology have expanded the application areas of deep learning methods, facilitating the transition from shallow to deep networks [10]. Researchers continue to develop optimized versions of DQL, such as updating neural network weights at each step to improve agent performance during training [7].
Reinforcement Learning: Reinforcement Learning (RL) is a branch of artificial intelligence that focuses on training agents to make decisions by interacting with environments to maximize cumulative rewards [11]. RL consists of fundamental components such as states, actions, policies, and reward signals [11]. Algorithms in RL are classified as model-based, model-free, value-based, and policy-based [11]. Sutton and Barto’s [12] seminal book provides a comprehensive guide covering the history, fundamental concepts, and algorithms of RL. The book consists of three parts: problem definition, basic solution methods, and advanced topics [12]. Recent developments in RL include the rise of deep reinforcement learning, which enhances RL’s capacity to solve complex, high-dimensional problems [11]. RL has found applications in various fields such as autonomous systems, healthcare, and energy management [11].
SOLID Principles: SOLID principles are an acronym combining the Single Responsibility Principle (SRP), Open-Closed Principle (OCP), Liskov Substitution Principle (LSP), Interface Segregation Principle (ISP), and Dependency Inversion Principle (DIP). The main objective of these principles is to increase software quality, reusability, scalability, and maintainability [13, 14].
II-A Network Topologies
An incorrect network topology can lead to problems such as performance degradation, data loss, high energy consumption, and critical system failures. Therefore, selecting the correct network topology is crucial [1]. Network topologies are generally divided into two types:
Physical Topology: Physical topology is the physical arrangement of connections and devices on the network (e.g., cable and device placements). Logical Topology: Logical topology describes the data movement scheme and communication paths within the network. Some popular Network Topology Types include:
Star Topology: In this network topology, all devices are arranged around a central hub. Each device connects with others through this central device. The central device regulates and controls data flow. It is frequently preferred due to its simplicity and ease of fault detection, but if the central device fails, the entire system goes down [15]. Dumbbell Topology: In this network topology, there are two ends, and these ends are connected to each other via a main link. Data between these ends is carried over the main link. The ends can behave like a local area network. It is used for traffic analysis and performance testing [2]. Mesh Topology: In this network topology, devices can establish multiple connections with each other. If every device is connected to every other device, it is called a full mesh; otherwise, it is called a partial mesh. It offers high-level data security and backup possibilities. Network resilience is high. Its disadvantages include high cost and complexity [15].
II-B TCP Protocol and Congestion Control
Developed in the 1970s, TCP/IP forms the foundation of modern internet communication [16]. TCP establishes virtual circuits between computers to manage data transmission as byte streams, while IP performs packet routing [17]. The protocol suite covers core network, transport, routing, and application layers [18]. TCP ensures reliable communication by segmenting large data groups, numbering packets, and using acknowledgment mechanisms to prevent data loss [17]. In 1983, TCP/IP replaced the Network Control Protocol as ARPAnet’s official data transport method [19]. Despite its utility, TCP/IP has security vulnerabilities such as IP spoofing and denial-of-service attacks [16]. To mitigate these risks, security enhancements like IPSec, VPNs, and encryption techniques have been developed [16, 19]. TCP/IP continues to evolve with innovations such as IP mobility, multicast, and IPv6, increasing its capacity and importance in modern networks [19]. The congestion window (cWnd) in TCP is an algorithm created to manage network congestion. This window detects the maximum number of packets a TCP flow can have in the network at any given time, directly affecting data transfer rate and network stability [20]. This concept, developed to prevent congestion collapse, ensures network functionality, especially in dense network traffic. Congestion collapse occurs when excessive data traffic overloads network resources, leading to a severe drop in performance [21]. TCP typically interprets packet loss as a congestion-related problem; however, this can cause issues in wireless networks where losses occur due to interference, signal attenuation, or mobility. To overcome such difficulties, modern congestion control algorithms include mechanisms to adaptively adjust the congestion window (e.g., reducing window size during idle or limited sending periods). The initial size and evolution of the congestion window have been re-developed over the years to improve performance and efficiency. In 2010, Allman made specific recommendations for determining and evolving the initial congestion window size, emphasizing a balance between fast data transfer and network stability [22]. Furthermore, innovative methods in congestion control, such as mechanisms like Explicit Congestion Notification (ECN) and hybrid approaches combining loss-based and delay-based control, optimize performance in different network environments [23]. TCP congestion control algorithms are evolving to address new challenges in modern networks. In 5G networks, high user density, large traffic flows, and mmWave communications pose unique obstacles for TCP performance [24]. Traditional loss-based congestion control algorithms struggle with rapid channel oscillations in 5G, leading to buffer bloat and delays [25]. To overcome these issues, researchers are developing dynamic congestion control mechanisms that adapt to network conditions. For example, D-TCP uses an Adaptive Increase/Adaptive Decrease approach based on available bandwidth [26], while FB-TCP uses fuzzy logic to adjust sending rates in urban 5G deployments [27]. These new algorithms aim to improve throughput, reduce packet loss and latency, and enhance fairness in various network environments. Machine learning techniques are also being investigated to optimize TCP congestion control in 5G networks [24]. The evolution of these methods over time ensures TCP remains compatible and effective in the constantly changing network landscape.
II-C Network Congestion and Performance Issues
Network congestion occurs when a network receives more data than it can process, leading to packet loss and unreliable connections. Congestion can result from hardware factors like aging routers, software issues like inefficient protocols, and other factors such as poor network design. Congestion causes a decrease in network performance, an increase in transmission delays, and a reduction in data transfer speed. As networks evolve to support multimedia applications, congestion continues to be a significant problem. Research has focused on understanding the causes of congestion and developing control solutions. Recent studies investigate neural networks for congestion control in various network environments. A deep reinforcement learning approach in wireless networks has shown significant improvements in throughput compared to traditional protocols [28]. For content-centric networks, a neural network-based algorithm has demonstrated an 85.53% increase in throughput by predicting and preventing queue overflow [29]. In wired networks, neural networks show promise in managing heterogeneous traffic flows and improving quality of service [30]. Additionally, neural networks have been applied to automatically adapt new users to existing quality of service policies in backbone networks and enhance traffic management efficiency [31]. These studies collectively demonstrate the potential of neural networks to address congestion control challenges in various network types, offering improved performance and adaptability compared to traditional methods.
II-D Deep Q-Learning
Deep Q-Learning (DQL) combines Q-learning with deep neural networks to address complex reinforcement learning problems with large state spaces [32]. While DQL has shown remarkable success, especially in Atari games [33], it suffers from overestimation of action values in certain situations [34, 33]. To mitigate this issue, researchers have proposed various solutions, including Double Q-learning, which has demonstrated improved performance in various games [33]. Furthermore, dropout techniques have been investigated as a way to reduce variance and overestimation in DQL, showing significant improvements in stability and performance [35]. Despite these advancements, DQL still faces challenges in real-world applications due to factors such as high-dimensional sensory inputs and the need for efficient environmental representations [32]. Ongoing research aims to increase the applicability and effectiveness of DQL in solving complex reinforcement learning problems across various domains.
II-E Reinforcement Learning
Reinforcement Learning (RL) is a branch of artificial intelligence focused on training agents to make decisions by interacting with environments to maximize cumulative rewards [11]. RL consists of fundamental components such as states, actions, policies, and reward signals [11]. Algorithms in RL are categorized as model-based, model-free, value-based, and policy-based [11]. The groundbreaking book by Sutton and Barto [12] provides a comprehensive guide covering the history, fundamental concepts, and algorithms of RL. The book comprises three sections: problem definition, basic solution methods, and advanced topics [12]. Recent developments in RL include the rise of deep reinforcement learning, which enhances RL’s capacity to solve complex, high-dimensional problems [11]. RL has found applications in various fields such as autonomous systems, healthcare, and energy management [11].
II-F Previous Works and Existing Methods
II-F1 Classical TCP Algorithms (Reno, CUBIC, BBR)
Reno: Initially operates in ”Slow Start” mode. In this mode, the congestion window (cwnd) typically starts from a small value and then increases exponentially. When network congestion occurs, this growth rate slows down; the exponential increase becomes linear. In cases of packet loss, the lost packet is retransmitted, and the algorithm transitions to fast recovery mode. Reno’s congestion control was first defined by Jacobson, significantly impacting the algorithm’s performance [3, 36]. Limitations and Disadvantages: It always attributes packet loss to network congestion, which is not always a correct assumption. This can lead to Reno unnecessarily slowing down the network and causing performance degradation. Packet loss can also originate from physical network conditions (signal weakness or interference) [37, 38]. In high-latency networks, it cannot react fast enough due to the delay in slow start processes. This typically weakens network performance due to congestion and delay issues observed in remote networks [38]. It is weak in managing multiple packet losses. Reno is designed for single packet losses and experiences performance degradation with multiple losses [38]. In wireless networks, it may misinterpret packet losses caused by physical conditions, further slowing down the network [37].
CUBIC: Initially, the cwnd value increases up to its maximum value. As it approaches, growth slows down, and then increases faster once it has passed. In case of congestion, the window size is reduced, and after congestion clears, it attempts to return to the maximum value [39]. Disadvantages and Limitations: In low-latency networks, it can cause congestion by excessively increasing the window size [39]. It detects network congestion only through packet loss and is insensitive to other causes. For example, it cannot detect differences in network conditions due to bandwidth limitations or changes in network topology [39]. It has a complex calculation mechanism, which creates an additional system resource load. This can negatively impact efficiency, especially on devices with low processing power [40].
BBR: Initially increases the data transfer rate to measure bandwidth and RTT values. To reduce the excess load generated during this phase, the cwnd value is temporarily decreased. The network’s bandwidth is continuously measured, and the location of the bottleneck is identified. To measure and update the RTT value, the cwnd value is periodically decreased, thus preventing buffer bloat. BBR is considered a significant step towards more efficient traffic control [4]. Disadvantages and Limitations: Due to its continuous bandwidth measurement, it can be misleading in situations where conditions change constantly. Especially when there are sudden fluctuations in the network, the BBR algorithm may not yield accurate results [4]. Incorrect detection of bottleneck points can lead to unnecessary reduction in data transfer rate or congestion. This can cause unnecessary data slowdowns in the network [4].
II-F2 Machine Learning for TCP
Recent research has explored the potential of machine learning to improve TCP congestion control algorithms by addressing the limitations of traditional approaches like Reno and CUBIC. These traditional algorithms often struggle with dynamic network conditions due to their fixed rules and limited adaptability [41, 42]. Machine learning-based approaches have shown better performance in various network scenarios, offering improved throughput-latency balance and adaptability to changing environments [43, 44]. Reinforcement learning, in particular, has demonstrated promise in developing adaptive TCP algorithms that can learn optimal congestion control policies through interaction with the network [41, 44]. For example, QTCP, a reinforcement learning-based approach, achieved 59.5% higher throughput while maintaining low latency compared to traditional TCP [44]. These developments indicate that machine learning can provide more dynamic and adaptive congestion control algorithms, potentially overcoming the limitations of traditional methods in complex and evolving network environments [42]. On the other hand, machine learning techniques, especially RL, can easily adapt to different conditions and continuously improve their learning, thereby increasing efficiency. Machine learning techniques can analyze complex data and learn from past experiences to make informed decisions about network congestion [45]. This enables data transmission with less loss and more efficient use of network capacity.
II-F3 RL-Based TCP Approaches
Recent studies have investigated various reinforcement learning (RL) approaches to improve TCP congestion control in diverse network environments. TCP-RLACC provides better data transmission and packet delivery in wireless networks by specifically adjusting the congestion window’s growth rate to network conditions [46]. TCP-Drinc utilizes deep reinforcement learning (Deep RL) to adjust the congestion window size, showing effective results in complex, dynamic networks [47]. A transfer learning-supported RL framework designed to reduce latency and increase data transfer rate in video streaming applications can distinguish between congestion-related and non-congestion-related losses [48]. Furthermore, a multi-agent RL approach used for TCP congestion control in multi-user networks performs better than traditional methods [48]. An RL-based TCP congestion avoidance mechanism in data center environments significantly reduces the impact of congestion by increasing end-to-end data transfer rate [49]. These studies clearly demonstrate the potential of RL-based approaches to improve TCP performance in various network scenarios.
II-G Application Areas
TCP congestion control algorithms are traditionally used to communicate between two computers, but their application areas are not limited to these. The Internet of Things (IoT) poses challenges for reliable communication among various devices. TCP, a reliable transport protocol, needs to be adapted for IoT environments. Various studies propose solutions to improve TCP performance in IoT contexts. Verma and Kumar [50] present a congestion control policy that rapidly adjusts transmission rates according to network conditions, enhancing throughput and fairness. Hamrioui and Lorenz [51] introduce LBA-Ie, a load balancing algorithm that integrates IoT parameters into TCP’s flow control, improving QoS and energy efficiency in e-health applications. Alam [52] proposes a framework addressing reliability issues in IoT device communication, focusing on error detection, correction, and acknowledgment at the TCP layer. Gomez et al. [53] provide guidance on implementing lightweight TCP stacks for constrained node networks, discussing simplification techniques and their tradeoffs. Collectively, these studies aim to adapt TCP to the unique challenges of IoT environments, enhancing reliability, efficiency, and performance. Vehicular communication technologies, including Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), and Vehicle-to-Everything (V2X), aim to enhance road safety by enabling information exchange between vehicles and infrastructure [54]. TCP plays a crucial role in reliable and secure data transmission for healthcare applications. It enables efficient transfer of patient data, medical images, and real-time videos for telemedicine [55]. TCP’s congestion control mechanism is vital for managing network traffic, and advanced versions like TCP CUBIC show improved performance in healthcare networks [56]. The increasing reliance on remote healthcare services necessitates robust network infrastructure and data management solutions, including big data and cloud computing [57].
II-H Research Gap and Study Contributions
As can be seen from the literature review, traditional TCP algorithms struggle to adapt to variable network conditions. All algorithms have their own advantages and disadvantages. Utilizing Reinforcement Learning’s Deep Q-networks model to combine these advantages into one algorithm and easily adapt to changing conditions is an innovative idea for the literature. This study, conducted in an area where similar works have not yet been performed in our country, will guide future studies and contribute to their further development.
III Methodology
III-A Overview
In this study, a Deep Q-Networks based Reinforcement Learning algorithm was developed to optimize TCP congestion control. Deep Q-Networks combine Q-learning, a Reinforcement Learning algorithm, with deep learning, enabling more optimal results in complex situations. The aim of this study was to improve network performance by real-time monitoring of network status and optimizing the cWnd value using Deep Q-Networks. The RL-based TCP algorithm was developed based on Markov Decision Processes. The Tensorflow library, Python, and C++ programming languages were used for the artificial intelligence processes during the algorithm’s development. This Reinforcement Learning-based TCP congestion control algorithm was tested in the Ns3 network simulation. The Ns3 simulation tool was run on the Linux operating system via the Windows Subsystem for Linux.
III-B Network Topology Setup
A network simulation was created to test and compare the analysis results of the developed TCP congestion control algorithm. ”Network Simulator 3” (Ns3) was used for the network simulation processes. The reason for choosing this network simulation is Ns3’s customizable structure and wide academic acceptance. NS3’s OpenGym module was used for Reinforcement Learning operations. This module integrates the simulation environment with an OpenAI Gym interface, allowing RL algorithms to be run on NS3.
Figure 1: Dumbbell Network Topology Visual
First, a Dumbbell network topology was created in the Ns3 simulation (Figure 1). In this created topology, there is 1 source node and 1 target node. The source and target nodes are connected to ”router-0” and ”router-1” devices with 10 Mbps links. The bottleneck link between ”router-0” and ”router-1” was configured as 2 Mbps. This transmission value is used to examine how traffic behaves in network conditions that respond quickly. The simulation duration is 10 seconds.
III-C RL-Based TCP Algorithm Design
III-C1 State Space
The current state of the network has been defined with features that enable the AI model to monitor existing network conditions and make appropriate decisions. The state space includes the following parameters that directly affect network performance:
  • BytesInFlight: The amount of data currently sent in the network, but not yet recognized.
  • Congestion Window (cWnd): The size of the congestion control window.
  • Round Trip Time (RTT): The time between sending a packet and receiving a response.
  • SegmentsAcked: The number of acknowledged segments.
III-C2 Action Space
The actions that the artificial intelligence model can perform have been defined in accordance with the congestion control mechanism of the TCP protocol. These actions are:
  • Increasing the Congestion Window: When the ”IncreaseWindow” function is called, the artificial intelligence model determines how to increase the cWnd value.
  • Decreasing the Congestion Window: The artificial intelligence model determines how much to decrease the cWnd value with the GetSsThresh function in case of congestion.
  • Keeping the Congestion Window Constant: When the IncreaseWindow or GetSsThresh functions are not called, the artificial intelligence model keeps the cWnd value constant.
These three basic actions allow the RL algorithm to dynamically adjust congestion control to optimize network performance.
III-C3 Reward Function
The reward function plays a critical role in guiding the Reinforcement Learning agent to achieve optimal network performance. In this study, a reward function was designed to encourage high throughput and low latency, reflecting the primary objectives of TCP congestion control. The reward function is defined as follows:
$\text{Reward}=\alpha\times\text{Throughput}-\beta\times\text{Latency}$
Where:
  • Throughput: Measured in Mbps, representing the data transfer rate.
  • Latency: Measured in milliseconds, representing the round-trip time.
  • $\alpha$ and $\beta$: Hyperparameters that determine the relative importance of throughput and latency. These values are tuned empirically to achieve the desired balance between the two performance metrics.
A positive reward is given for an increase in throughput and a decrease in latency, while a penalty is imposed for the opposite. This structure encourages the agent to learn policies that maximize data transfer efficiency while minimizing delays.
III-C4 Training and Decision Process
The Deep Q-Networks (DQN) model was trained using the collected state and reward data from the NS3 simulation environment. The training process involved iteratively updating the Q-network’s weights based on the Bellman equation, aiming to minimize the temporal difference (TD) error. The training process can be summarized in the following steps:
  • Initialization: The Q-network is initialized with random weights. An experience replay buffer is also initialized to store past (state, action, reward, next_state) transitions.
  • Exploration vs. Exploitation: During training, the agent uses an $\epsilon$-greedy policy to balance exploration (trying new actions) and exploitation (choosing the best-known action). The $\epsilon$ value gradually decays over time, encouraging more exploitation as the agent learns.
  • Interaction with Environment: At each time step in the NS3 simulation, the agent observes the current network state (BytesInFlight, cWnd, RTT, SegmentsAcked).
  • Action Selection: Based on the observed state and the $\epsilon$-greedy policy, the agent selects an action (IncreaseWindow, DecreaseWindow, Keep Constant).
  • Execution and Observation: The selected action is executed in the NS3 environment, leading to a new network state and a corresponding reward.
  • Experience Storage: The transition (current_state, action, reward, next_state) is stored in the experience replay buffer.
  • Batch Sampling and Training: Periodically, a mini-batch of experiences is randomly sampled from the replay buffer. These samples are used to train the Q-network, updating its weights to predict Q-values more accurately. A separate target Q-network, updated less frequently, is used to stabilize the training process.
  • Iteration: Steps 3-7 are repeated for a predefined number of training epochs or until the agent’s performance converges.
The decision-making process in the deployed RL-based TCP algorithm is continuous. In real-time, the model observes the current network state, feeds this information into its trained Q-network, and selects the action that maximizes the predicted Q-value for that state. This allows for dynamic and adaptive congestion control in response to changing network conditions.
IV Experimental Results
The experimental evaluation of the Deep Reinforcement Learning-based TCP congestion control algorithm in the Ns3 simulation environment yielded significant findings regarding its performance compared to traditional methods, specifically TCP New Reno.
Latency Improvement: The RL-based algorithm demonstrated a notable reduction in network latency. On average, the proposed algorithm achieved 12.51% lower latency compared to TCP New Reno in various simulated network conditions. This indicates the algorithm’s ability to minimize packet delays, which is crucial for real-time applications. The reduction in Round-Trip Time (RTT) is clearly visible in Figure 2.
Figure 2: Comparison of Round-Trip Time (RTT) for Deep Reinforcement Learning-based TCP Algorithm and TCP New Reno.
Throughput Enhancement: The data transmission rate (throughput) was significantly higher with the RL-based algorithm. An average increase of 68.31% in throughput was observed compared to TCP New Reno, highlighting the algorithm’s effectiveness in utilizing network capacity more efficiently and maximizing data flow, as depicted in Figure 3.
Figure 3: Comparison of Average Throughput for Deep Reinforcement Learning-based TCP Algorithm and TCP New Reno.
Adaptability to Variable Network Conditions: During simulation tests, the RL-based algorithm exhibited strong adaptability to dynamic changes in network conditions (e.g., sudden traffic fluctuations, varying bottleneck bandwidth). The algorithm was able to quickly adjust its cWnd value to optimize performance, demonstrating its robustness in unpredictable environments, unlike traditional algorithms that rely on fixed rules. This adaptive behavior of the congestion window can be observed in Figure 4.
Figure 4: Comparison of Congestion Window (cWnd) for Deep Reinforcement Learning-based TCP Algorithm and TCP New Reno.
Figure 5: Reward Change During Training for the Deep Reinforcement Learning-based TCP Algorithm.
These findings support the hypothesis that a Reinforcement Learning-based TCP congestion control algorithm can provide higher network throughput, lower latency, and faster adaptation to network conditions compared to traditional algorithms.
V Conclusion and Future Work
This study successfully developed and evaluated an innovative TCP congestion control algorithm based on Deep Reinforcement Learning, specifically using Deep Q-Networks. The primary objective was to address the limitations of traditional TCP algorithms in adapting to dynamic and variable network conditions, which often result in suboptimal performance metrics such as high latency and low throughput. The experimental results obtained from the Ns3 simulation environment strongly support the effectiveness of the proposed RL-based algorithm. It consistently outperformed TCP New Reno, a widely used traditional algorithm, by achieving an average of 12.51% lower latency and an impressive 68.31% higher data transmission rate. These quantitative improvements underscore the algorithm’s ability to significantly enhance network performance. Furthermore, the simulations demonstrated the algorithm’s superior adaptability to fluctuating network conditions, a critical advantage over static, rule-based approaches. This adaptability suggests that the AI-based algorithm can maintain high efficiency even in unpredictable network environments. The successful application of Markov Decision Processes and Deep Q-Networks confirms that this machine learning paradigm can provide a correct and efficient solution for complex TCP congestion control problems. The use of TensorFlow, Python, and C++ proved to be effective in developing and simulating an intelligent agent. The implications of these findings are particularly significant for modern network infrastructures, especially mobile networks and IoT devices, where network conditions are inherently variable and resource constraints are common. By dynamically optimizing the cWnd value, the algorithm can contribute to more stable and efficient data communication in these challenging environments. Although the results are promising, it is important to acknowledge certain considerations. The performance of the RL algorithm is highly dependent on the design of the reward function and the quality of the training data. Future research could explore alternative reward structures and more complex network topologies to further validate and generalize the algorithm’s performance. In conclusion, this study provides compelling evidence that Deep Reinforcement Learning offers a powerful paradigm for developing highly adaptive and efficient TCP congestion control algorithms. The developed solution contributes significantly to the field of network management by offering a robust approach to optimizing network performance in increasingly dynamic and complex communication landscapes.
VI Recommendations
Based on the findings and conclusions of this study, the following recommendations are proposed for future research and development.
  • Exploration of More Complex Reward Functions: Investigate alternative and more sophisticated reward functions that might incorporate additional network metrics (e.g. fairness, packet loss rate as a direct penalty, energy efficiency) to guide the RL agent towards more holistic optimization.
  • Testing in Diverse Network Topologies: Expand the evaluation of the algorithm to more complex and realistic network topologies beyond the dumbbell model, such as fat-tree topologies common in data centers or mesh networks found in IoT environments, to assess its scalability and generalizability.
  • Integration with Real-World Network Testbeds: Transition from simulation-based testing to deployment and evaluation on actual network testbeds. This would provide more realistic insights into the algorithm’s performance under real-world traffic patterns, hardware limitations, and unforeseen network events.
  • Comparative Analysis with Other Advanced TCP Variants: Conduct a more extensive comparative analysis including other advanced TCP congestion control algorithms (e.g., BBR, Vegas, Hybla) and recent machine learning-based approaches from the literature to precisely quantify the proposed algorithm’s advantages.
  • Hyperparameter Optimization Techniques: Employ advanced hyperparameter optimization techniques (e.g., Bayesian optimization, genetic algorithms) to fine-tune the RL model’s parameters (e.g., learning rate, discount factor, exploration rate decay) for even better performance.
  • Transfer Learning and Meta-Learning: Explore the application of transfer learning or meta-learning techniques to enable the RL agent to quickly adapt to entirely new network environments or sudden, drastic changes without requiring extensive re-training from scratch.
  • Security and Robustness Analysis: Conduct a thorough analysis of the algorithm’s robustness against potential adversarial attacks or unexpected network behaviors, ensuring its security and reliability in production environments.
  • Resource Footprint Analysis: Evaluate the computational and memory footprint of the RL-based algorithm, especially for deployment on resource-constrained devices like IoT nodes, and explore optimization strategies for lightweight implementation.